OpenStudy (johnnewman):

The length of one leg of a right triangle is 21 ft and the length of the hypotenuse is 29 ft. What is the length of the other leg? A. 50.0 feet B. 25.0 feet C. 20.0 feet D. 35.8 feet

OpenStudy (mathstudent55):

Are you familiar with the Pythagorean theorem?

OpenStudy (johnnewman):

yes I am a^2 + b^2 = c^2

OpenStudy (johnnewman):

so 21^2 + 29^2 = c^2 right

OpenStudy (johnnewman):



OpenStudy (mathstudent55):

No. You need to be careful. You have the Pythagorean equation correct, but you need to use a, b, and c with the correct sides. a and b are always the legs, the sides that form the right angle. c is always the hypotenuse, the side opposite the right angle and always the longest side.

OpenStudy (mathstudent55):